# Modulo Sampling: Performance Guarantees in The Presence of Quantization

### v1

@@ Abstract
In this paper, we investigate the relationship between the dynamic range and quantization noise power in modulo analog-to-digital converters (ADCs). Two modulo ADC systems are considered: (1) a modulo ADC which outputs the folded samples and an additional 1-bit folding information signal, and (2) a modulo ADC without the 1-bit information. A recovery algorithm that unfolds the quantized modulo samples using the extra 1-bit folding information is analyzed. Using the dithered quantization framework, we show that an oversampling factor of $\mathrm{OF}>3$ andt a quantizer resolution of $b>3$ are sufficient conditions to unfold the modulo samples. When these conditions are met, we demonstrate that the mean squared error (MSE) performance of modulo ADC with an extra 1-bit folding information signal is better than that of a conventional ADC with the same number of bits used for amplitude quantization. Since folding information is typically not available in modulo ADCs, we also propose and analyze a recovery algorithm based on orthogonal matching pursuit (OMP) that does not require the 1-bit folding information. In this case, we prove that $\mathrm{OF}>3$ and $b>3+\log_{2}(\delta)$ for some $\delta>1$ are sufficient conditions to unfold the modulo samples. For the two systems considered, we show that, with sufficient number of bits for amplitude quantization, the mean squared error (MSE) of a modulo ADC is $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ whereas that of a conventional ADC is only $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$. We extend the analysis to the case of simultaneous acquisition of weak and strong signals occupying different frequency bands. Finally, numerical results are presented to validate the derived performance guarantees.
@@ I Introduction
Most modern data acquisition systems are equipped with analog-to-digital converters (ADCs) to transform continuous-time observations to some format suitable for processing in the digital domain. Standard ADCs first sample the signal at equally-spaced time intervals and then the amplitudes of the samples are mapped to a finite set of discrete values by a quantizer [1]. The aim is to accurately represent the original analog signal in the digital domain. One critical consideration in an ADC design is ensuring that the dynamic range (DR) of the ADC exceeds that of the input signal, where DR refers to the difference between the smallest and largest values that can be represented. When the DR of the ADC falls short of capturing the full range of the input signal, clipping occurs [2, 3, 4]. This loss of information compromises the accuracy of the digital representation and impacts the reliability of subsequent data processing.
@@
One of the main focuses of existing works on modulo ADC is to develop a robust and computationally efficient recovery algorithm that operates close to the Nyquist rate. The existing literature presents diverse recovery algorithm strategies. These encompass higher-order-differences-based approach [9], prediction-based method [10], wavelet-based technique [11], Fourier-domain methodologies [12, 13, 14, 15], and others. Among these methods, [13, 14, 15, 16] are identified to be robust recovery algorithms approaching the Nyquist rate. Apart from recovery algorithms, there are also efforts to advance the hardware implementation of modulo ADCs [12, 17, 18, 16, 19, 20]. Furthermore, USF has been studied extensively across various signal models, such as Finite-Rate-of-Innovation (FRI) signals [21], sparse signals [22, 23, 24], graph signals [25], and shift-invariant spaces [26, 27], among others.
Nevertheless, research addressing the impact of quantization on modulo ADCs is notably limited. In [9], the authors introduced a higher-order differences-based recovery algorithm for bandlimited (BL) signals that operates at approximately $2\pi e\;(\approx 17)$ times the Nyquist rate when there is no quantization. Theoretical guarantees for the higher-order show that this sufficient condition for the sampling rate becomes stricter in the presence of bounded noise [9]. To improve upon this, a prediction filter-based algorithm to unfold the modulo ADC output was proposed [10]. When the prediction filter is sufficiently long, this approach enables signal recovery at a rate close to the Nyquist rate. However, this recovery method fails in the presence of quantization noise. Recent work on USF radars [28] has investigated the impact of quantization on a modulo ADC if the input signal contains weak and strong components, however, this work uses higher-order differences to unfold the modulo samples. Signal-to-quantization noise ratio (SQNR) of modulo ADCs has been analyzed in [29] but the reconstruction method used relies on a separately encoded 2-bit reset signal.
This work aims to address this gap and provide a more comprehensive understanding of how modulo sampling and quantization interact, particularly in scenarios involving BL signals with diverse target strengths. Moreover, this study identifies suitable conditions for which a modulo ADC has better performance than a conventional ADC without modulo when the number of bits used for amplitude quantization are the same. The following are major contributions:
@@
    We analyze the recovery algorithm introduced in [16], which leverages the bits distribution mechanism to unfold the modulo ADC output. In [16], the algorithm allocates 1-bit from the total bit budget to detect modulo folding events. In contrast, this work assumes the full ADC resolution ($b$) is used for quantization, with an additional 1-bit folding information explicitly provided by the modulo ADC. Within a dithered quantization framework, we evaluate the reconstruction performance of this approach. We establish sufficient conditions on the oversampling factor ($\mathrm{OF}$) and $b$ to ensure that the modulo ADC offers superior quantization noise suppression compared to conventional ADCs without the modulo operator. Specifically, we prove that $\mathrm{OF}>3$ and $b>3$ are enough for the modulo ADC to achieve lower mean squared error ($\mathrm{MSE}$) than conventional ADCs.
@@
    Since the 1-bit folding information is generally unavailable in the modulo ADC, we propose and examine an orthogonal matching pursuit (OMP)-based recovery algorithm to unfold the modulo ADC output. Our analysis demonstrates that ensuring $\mathrm{OF}>3$ and $b>3+\log_{2}(\delta)$, where $\delta>1$ is some constant, is sufficient for achieving superior quantization noise suppression in modulo ADCs over conventional ADCs. Additionally, we present a case study that analyzes a BL signal having both weak and strong components occupying different frequency bands and demonstrate that a modulo ADC outperforms a conventional ADC in capturing the strong and weak signal components simultaneously.
@@
    We show that for the aforementioned settings, the MSE of the modulo ADC is on the order of $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$, while that of a standard ADC is on the order of $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$. This underscores the importance of oversampling in modulo ADC in retrieving the original signal from quantized modulo observations. Finally, we validate the proposed theory with numerical simulations.
The paper is organized as follows: The two modulo ADC system models considered in the study and our proposed recovery methods are described in Section II. Theoretical MSE performance guarantees for the recovery algorithms are established in Section III. Then, we specialize our derived performance guarantees to the case of simultaneous acquisition of weak and strong signals in Section IV. Numerical results are provided in Section V to validate our proposed theory. Finally, we summarize the work in Section VI.
Notation: The following notations are used throughout the paper. We denote the set of real numbers, integers, and natural numbers as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{N}$, respectively. The space of square-integrable functions on $\mathbb{R}$ is represented by $L^{2}(\mathbb{R})$, and $\mathbb{E}[\cdot]$ denotes the expectation operator. We use bold lowercase letters to represent vectors, and bold capital letters to denote matrices. The $\ell_{p}$-norm of a vector $\mathbf{z}$ is denoted as $\|\mathbf{z}\|_{p}$, with a similar notation extending to $\ell_{p}$-norms of functions. When referring to discrete-time signals, we employ the notation $z[n]$ to signify $z(nT_{\mathrm{s}})$, assuming the sampling period $T_{\mathrm{s}}$ is evident from the context. For a discrete-time signal $x[n]$, the first-order difference operator is defined as $\Delta x[n]=x[n]-x[n-1]$, under the assumption that $x[-1]=0$. To describe the growth rate of a quantity concerning some parameter, we employ standard asymptotic notation, such as $\mathcal{O}(\cdot)$. The cumulative sum operation applied to a vector $\mathbf{x}\in\mathbb{R}^{N\times 1}$ yields a vector $\mathbf{y}\in\mathbb{R}^{N\times 1}$, where the $k^{\text{th}}$ element of $\mathbf{y}$ is given by $y_{k}=\sum_{i=1}^{k}x_{i}$. The pseudo-inverse of a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, where $n\gg m$, is denoted by $\mathbf{A}^{\dagger}$, and defined as $\mathbf{A}^{\dagger}=(\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}^{T}$.
@@ II-A Modulo ADC System With Extra 1-Bit Folding Information
We consider the modulo ADC system shown in Fig. 2a. The input $f(t)\in L^{1}(\mathbb{R})$ is a bandlimited signal with frequency support $\left[-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right]$. The input signal is fed to a non-linear modulo sampling block, denoted $\mathcal{M}_{\lambda^{\prime}}(\cdot)$, to produce the discrete-time modulo observations $f_{\lambda^{\prime}}[n],\;n\in\mathbb{Z}$. Mathematically,
@@
[…]
where $\lambda^{\prime}\in(0,\|f(t)\|_{\infty})$ is the modulo operator threshold. The range of $\lambda^{\prime}$ is chosen to exclude the trivial case that the input signal amplitude is within the linear region of the modulo. The sampling rate is $\omega_{s}=\frac{2\pi}{T_{s}}=\mathrm{OF}\times\omega_{m}$, where $\mathrm{OF}>1$. We will also use $\rho=\frac{1}{\mathrm{OF}}$ in the analysis.
The folded samples are fed to a $b$-bit scalar quantizer $\mathcal{Q}_{b}(\cdot)$ with dynamic range $[-\lambda,+\lambda]$. The nonlinearity of the quantization process makes it difficult to analyze the impact of quantization in modulo sampling. To this end, a dither signal $d[n]$ is added to $f_{\lambda^{\prime}}[n]$ prior to quantization. An independent and identically-distributed (i.i.d.) triangle noise sequence with amplitude support $\big{(}-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\big{]}$ is used for the dither signal $d[n]$. The rationale for incorporating a triangle dither in the system model is twofold. First, the modulo operation can be set so that $f_{\lambda^{\prime}}[n]+d[n]$ will not overload the quantizer. This, together with the properties of triangle dither, guarantees that Shuchman conditions [30] are satisfied, which allows the first-order and second-order statistics of the quantization noise to be derived from the system parameters. Second, the sequences $\epsilon[n]$ and $f_{\lambda^{\prime}}[n]$ become uncorrelated and the conditional second-order moment of $\epsilon[n]$ becomes independent of $f_{\lambda^{\prime}}[n]$, i.e., $\mathbb{E}\left[\epsilon^{2}[n]|f_{\lambda^{\prime}}[n]\right]=\mathbb{E}\left[\epsilon^{2}[n]\right]$.
Each quantization bin has a width of $\frac{2\lambda}{2^{b}}$. To prevent the signal from overloading the quantizer, we set the modulo threshold to be $\lambda^{\prime}=\frac{(2^{b}-2)\lambda}{2^{b}}$. The quantizer output can then be written as
@@
In addition to $f_{\lambda^{\prime}}[n]$, the modulo sampling block also generates a 1-bit discrete-time binary signal $c[n]$ which contains information about the $(2\mathbb{Z}+1)\lambda^{\prime}$ level crossings in $f(t)$. More precisely, $c[n]=1$ whenever the input signal crosses the level $(2\mathbb{Z}+1)\lambda^{\prime}$ within the time interval $(nT_{s},(n+1)T_{s}]$ while $c[n]=0$ otherwise. The set of indices $n$ where $c[n]=1$ is denoted as $\mathcal{S}$. There are several ways to generate this 1-bit folding information. One approach is explored in [16], which only incurs one additional OR gate to the setup. This penalty is fixed regardless of $b$. From a power efficiency viewpoint, this additional OR gate has minimal impact on the power consumption of the modulo ADC.

### v2

@@ Abstract
In this paper, we investigate the trade-off between dynamic range and quantization noise in modulo analog-to-digital converters (ADCs). We study two modulo ADC architectures: (1) a modulo ADC that outputs folded samples along with an additional 1-bit folding information signal, and (2) a modulo ADC that outputs only the folded samples. For the first system, we analyze a recovery algorithm that leverages the folding information to correctly unfold the quantized samples. Within the dithered quantization framework, we show that an oversampling factor $\mathrm{OF}>3$ and quantizer resolution $b>3$ are sufficient to guarantee the successful unfolding of the samples. Under these conditions, the system achieves a mean squared error (MSE) lower than that of a conventional ADC with the same number of amplitude quantization bits. Since folding information is typically unavailable in practical modulo ADCs, we also propose an orthogonal matching pursuit (OMP)-based recovery algorithm for the second system, which relies solely on the folded samples. In this case, we prove that the samples can be unfolded accurately if $\mathrm{OF}>3$ and $b>3+\log_{2}(\zeta)$, for some penalty term $\zeta>1$. For both systems, we show that the MSE of a modulo ADC is in $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$ when there is a sufficient number of bits for amplitude quantization. In contrast, the MSE of a conventional ADC is only in $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$. We further extend our analysis to the simultaneous acquisition of weak and strong signals occupying different frequency bands. Finally, numerical results are provided to validate the theoretical performance guarantees established.
@@ I Introduction
Most modern data acquisition systems use analog-to-digital converters (ADCs) to transform continuous-time observations into a format suitable for digital processing. Standard ADCs first sample the signal at equally-spaced time intervals and then map the amplitude samples to a finite discrete set of values using a quantizer [1, 2]. The aim is to accurately represent the original analog signal in the digital domain. One critical consideration in an ADC design is ensuring that the dynamic range (DR) of the ADC exceeds that of the input signal, where DR refers to the difference between the smallest and largest values that can be represented. When the DR of the ADC falls short of capturing the full range of the input signal, clipping occurs [3, 4, 5]. This loss of information compromises the accuracy of the digital representation and impacts the reliability of subsequent data processing. Moreover, obtaining a signal with lower DR can reduce the ADC’s power consumption and can play a vital role in energy-constrained systems [6].
@@
A key focus of existing research on modulo ADCs is developing robust, computationally efficient recovery algorithms that operate near the Nyquist rate. The existing literature presents diverse recovery algorithm strategies. These encompass higher-order-differences-based approach [12, 11], prediction-based method [13], wavelet-based technique [14], Fourier-domain methodologies [15, 16, 17, 18, 19], compressed sensing-based approaches [20, 21, 22], and others. Among these methods, [21, 15, 17, 20, 19] are identified to be robust recovery algorithms approaching the Nyquist rate. Apart from recovery algorithms, there are also efforts to advance the hardware implementation of modulo ADCs [16, 23, 24, 21, 25, 26, 27]. Furthermore, USF has been studied extensively across various signal models, such as Finite-Rate-of-Innovation (FRI) signals [28], sparse signals [29, 30, 31], graph signals [32], multi-tone signals [33], and shift-invariant spaces [34], among others.
Nevertheless, research addressing the impact of quantization on modulo ADCs is notably limited. In [12], the authors introduced a higher-order differences (HoD)-based recovery algorithm for bandlimited (BL) signals that operates at approximately $2\pi e\;(\approx 17)$ times the Nyquist rate when there is no quantization. Theoretical guarantees for the HoD-based recovery show that this sufficient condition for the sampling rate increases in the presence of bounded noise [11]. Recently, it was shown that this penalty in sampling rate under a bounded noise setting can be removed with high probability by using a multi-channel modulo sampler [35]. To improve upon the HOD approach [12], a prediction filter-based algorithm to unfold the modulo ADC output was proposed [13]. When the prediction filter is sufficiently long, this approach enables signal recovery at a rate close to the Nyquist rate. However, this recovery method fails in the presence of quantization noise. In [36], the authors proposed a thresholding-based recovery algorithm with performance guarantees under bounded noise. However, their approach is based on a modulo ADC model that includes hysteresis and folding transients to account for hardware non-idealities. The recovery error guarantees established in [36, Theorem 2] become trivial under zero hysteresis. Recent work on USF radars [37] has investigated the impact of quantization on a modulo ADC if the input signal contains weak and strong components, however, this work uses HoD-based recovery to unfold the modulo samples. In [38], the authors focus on signal-to-quantization-noise ratio (SQNR) for inputs with varying amplitude distributions and emphasize hardware implementation, using a reconstruction method that relies on an encoded 2-bit reset signal.
This work aims to address this gap and provide a more comprehensive understanding of how modulo sampling and quantization interact, particularly in scenarios involving BL signals with diverse target strengths. Moreover, this study identifies suitable conditions for which a modulo ADC has better performance than a conventional ADC without modulo when the number of bits used for amplitude quantization is the same. The following are the major contributions of our work:
@@
    We analyze the recovery algorithm [21] that uses the quantized modulo samples and 1-bit folding information signal to unfold the modulo ADC output. In this setup, the 1-bit folding information [^1] Note that the implementation in [21] uses one bit from the $b$-bit ADC to detect modulo folding events. In contrast, this work uses a $b$-bit ADC for amplitude quantization and an additional 1-bit ADC to indicate folding events. The $b$ in [21] needs to be increased by 1 to make the two modulo ADC implementations comparable in terms of bits per second. indicates the positions of the modulo samples where folding occurred. Within a dithered quantization framework [39, 40, 2], we evaluate the reconstruction performance of this approach. We establish sufficient conditions on the oversampling factor ($\mathrm{OF}$) and $b$ to ensure that the modulo ADC offers superior quantization noise suppression compared to conventional ADCs without the modulo operator. Specifically, we prove that $\mathrm{OF}>3$ and $b>3$ are sufficient for the modulo ADC to achieve lower mean squared error ($\mathrm{MSE}$) than conventional ADCs.
@@
    Since the 1-bit folding information is generally unavailable in the modulo ADC, we propose and examine an orthogonal matching pursuit (OMP)-based recovery algorithm to unfold the modulo ADC output. Our analysis demonstrates that having $\mathrm{OF}>3$ and $b>3+\log_{2}(\zeta)$, for some penalty term $\zeta>1$, are sufficient for the modulo ADC without 1-bit folding information to achieve the same MSE performance guarantees as those of the modulo ADC with 1-bit folding information. Additionally, we present a case study that analyzes a BL signal having both weak and strong components occupying different frequency bands and demonstrate that a modulo ADC outperforms a conventional ADC in capturing the strong and weak signal components simultaneously.
@@
    We show that, for the aforementioned settings, the MSE of the modulo ADC is on the order of $\mathcal{O}\left(\frac{1}{\mathrm{OF}^{3}}\right)$, while that of a standard ADC is on the order of $\mathcal{O}\left(\frac{1}{\mathrm{OF}}\right)$. This underscores the importance of oversampling in retrieving the original signal from quantized modulo observations. We validate the proposed theory with numerical simulations.
A portion of the aforementioned contributions was previously presented in [41], where a modulo ADC with explicit 1-bit folding information was analyzed within a dithered quantization framework. Building on this prior work, the present paper addresses the practical issue of modulo ADC without the 1-bit folding information by introducing an OMP-based signal recovery and establishing its MSE performance guarantees. Additionally, we extend the analysis to BL signals with both strong and weak components.
The paper is organized as follows: The two modulo ADC system models considered in the study and our proposed recovery methods are described in Section II. Theoretical MSE performance guarantees and computational complexities of the recovery algorithms are established in Section III. Then, we specialize our derived performance guarantees to the case of simultaneous acquisition of weak and strong signals in Section IV. Numerical results are provided in Section V to validate our proposed theory. Finally, we summarize the work in Section VI.
Notation: The following notations are used throughout the paper. We denote the set of real numbers, integers, and natural numbers as $\mathbb{R}$, $\mathbb{Z}$, and $\mathbb{N}$, respectively. The space of square-integrable functions on $\mathbb{R}$ is represented by $L^{2}(\mathbb{R})$, and $\mathbb{E}[\cdot]$ denotes the expectation operator. We use bold lowercase letters to represent vectors, and bold capital letters to denote matrices. The $\ell_{p}$-norm of a vector $\mathbf{z}$ is denoted as $\|\mathbf{z}\|_{p}$, with a similar notation extending to $\ell_{p}$-norms of functions. When referring to discrete-time signals, we employ the notation $z[n]$ to signify $z(nT_{\mathrm{s}})$, assuming the sampling period $T_{\mathrm{s}}$ is evident from the context. For a discrete-time signal $x[n]$, the first-order difference operator is defined as $\Delta x[n]=x[n]-x[n-1]$, under the assumption that x[1]=0x[-1]=0. To describe the growth rate of a quantity concerning some parameter, we employ standard asymptotic notation, such as $\mathcal{O}(\cdot)$. The cumulative sum operation applied to a vector $\mathbf{x}\in\mathbb{R}^{N\times 1}$ yields a vector $\mathbf{y}\in\mathbb{R}^{N\times 1}$, where the $k^{\text{th}}$ element of $\mathbf{y}$ is given by $y_{k}=\sum_{i=1}^{k}x_{i}$. The pseudo-inverse of a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, where $n\gg m$, is denoted by $\mathbf{A}^{\dagger}$, and defined as $\mathbf{A}^{\dagger}=(\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}^{T}$. We used $\omega$ and $\Omega$ to denote the frequencies for the CTFT and DTFT spectrum, respectively.
@@ II-A Modulo ADC System With Extra 1-Bit Folding Information
We consider the modulo ADC system shown in Fig. 2a. The input $f(t)\in L^{1}(\mathbb{R})$ is a bandlimited signal with frequency support $\left(-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right)$. The input signal is fed to a non-linear modulo sampling block, denoted $\mathcal{M}_{\lambda^{\prime}}(\cdot)$, to produce the discrete-time modulo observations $f_{\lambda^{\prime}}[n],\;n\in\mathbb{Z}$. Mathematically,
@@
$=[(f(nT_{s})+\lambda^{\prime})\;\mathrm{mod}\;2\lambda^{\prime}]-\lambda^{\prime},$ (1)
where $\lambda^{\prime}\in(0,\|f(t)\|_{\infty})$ is the modulo operator threshold. The range of $\lambda^{\prime}$ is chosen to exclude the trivial case that the input signal amplitude is within the linear region of the modulo. The sampling rate is $\omega_{s}=\frac{2\pi}{T_{s}}=\mathrm{OF}\times\omega_{m}$, where $\mathrm{OF}>1$. We define $\rho=\frac{1}{\mathrm{OF}}$ for our analysis. Effectively, the frequency support of $f[n]=f(nT_{\mathrm{s})}$ is $(-\rho\pi,+\rho\pi)$.
The folded samples are fed to a $b$-bit scalar quantizer $\mathcal{Q}_{b}(\cdot)$ with dynamic range $\lambda,+\lambda]$. The nonlinearity of the quantization process makes it difficult to analyze the impact of quantization in modulo sampling. To this end, a dither signal $d[n]$ is added to $f_{\lambda^{\prime}}[n]$ prior to quantization. An independent and identically-distributed (i.i.d.) triangle noise sequence with amplitude support $\big{(}-\frac{2\lambda}{2^{b}},+\frac{2\lambda}{2^{b}}\big{]}$ is used for the dither signal $d[n]$. The rationale for incorporating a triangle dither in the system model is twofold. First, the modulo operation can be set so that $f_{\lambda^{\prime}}[n]+d[n]$ will not overload the quantizer. This, together with the properties of triangle dither, guarantees that Shuchman conditions [39] are satisfied, which allows the first-order and second-order statistics of the quantization noise to be derived from the system parameters. Second, using [40, Theorem 2], the sequences $\epsilon[n]$ and $f_{\lambda^{\prime}}[n]$ become uncorrelated and the conditional second-order moment of $\epsilon[n]$ becomes independent of $f_{\lambda^{\prime}}[n]$, i.e., $\mathbb{E}\left[\epsilon^{2}[n]|f_{\lambda^{\prime}}[n]\right]=\mathbb{E}\left[\epsilon^{2}[n]\right]$.
Each quantization bin has a width of $\frac{2\lambda}{2^{b}}$. To prevent the quantizer from being overloaded by the folded signal, we set the ADC DR to be $\lambda=\frac{2^{b}\lambda^{\prime}}{(2^{b}-2)}$. The quantizer output can then be written as
@@
In addition to $f_{\lambda^{\prime}}[n]$, the modulo sampling block also generates a 1-bit discrete-time binary signal $c[n]$ which contains information about the $(2\mathbb{Z}+1)\lambda^{\prime}$ level crossings in $f(t)$. More precisely, $c[n]=1$ whenever the input signal crosses the level $(2\mathbb{Z}+1)\lambda^{\prime}$ within the time interval $(nT_{s},(n+1)T_{s}]$ while $c[n]=0$ otherwise. The set of indices $n$ where $c[n]=1$ is denoted as $\mathcal{S}$.
There are several ways to generate this 1-bit folding information. One approach is explored in [21], which only incurs one additional OR gate to the setup. This penalty is fixed regardless of $b$. From a power efficiency viewpoint, this additional OR gate has minimal impact on the power consumption of the modulo ADC. In FPGA implementations, the OR gate can be realized using a small portion of a look-up table (LUT), resulting in negligible resource usage and power overhead. In ASIC implementations, it can be built using standard CMOS logic, typically requiring only a few transistors with minimal area and energy consumption. Since these implementation paths are straightforward and well-supported by existing digital design flows, the prototype in [21] remains practical and scalable, even when extended to higher-bit resolutions. However, we note that challenges may arise in very high-speed applications where generating this 1-bit side information and synchronizing it with the folded samples become more difficult.